# Analyzing the Influence of Geometrical Deformation on Photon Sphere and Shadow Radius: A New Analytical Approach - Part I: Spherically Symmetric Spacetimes

### v2

@@ Abstract
In this paper, we present a novel approach to investigate the behavior of the photon sphere and shadow radius. Our method leverages extended gravitational decoupling and unveils two key analytic results. First, the additional matter field alters the photon sphere radius: it increases if $g^{\prime}(r_{ph}^{(0)})>0$ and decreases if $g^{\prime}(r_{ph}^{(0)})<0$ (where $g^{\prime}$ represents the derivative of a specific metric function evaluated at the original photon sphere radius). Second, the presence of the matter field can modify the black hole shadow size. If $g\left(r_{ph}^{(0)}\right)>0$, the shadow shrinks, while it grows for $g\left(r_{ph}^{(0)}\right)<0$. These findings offer a deeper understanding of how matter distribution affects the characteristics of black holes and their observable features. By providing a systematic framework supported by diverse illustrative examples, our investigation not only sheds light on these fundamental aspects but also quantifiably advances the theoretical framework in black hole astrophysics.
@@ I Introduction
One of the most compelling experimental pieces of evidence supporting the existence of black holes is provided by the images capturing the shadow of a black hole, as obtained by the Event Horizon Telescope Collaboration bib:eht1 ; bib:eht2 ; bib:eht3 . The pioneering work on light deflection by gravitationally intense stars is attributed to Synge bib:singe . Bardeen calculated the radius of a photon sphere for a Schwarzschild black hole, demonstrating that its radius equals $3M$ bib:bardeen1 . This photon sphere manifests as a shadow, appearing as a dark spot in the observer’s sky, with a radius of $3\sqrt{3}M$ (roughly 2.5 times larger than the radius of the event horizon, $r=2M$). He also demonstrated that in the case of a rotating object, the shape of the shadow undergoes changes. However, in reality, astrophysical black holes are not isolated objects; they exist within an expanding universe and are often surrounded by luminous accretion flows, factors that significantly influence our observations. Therefore, to accurately describe a black hole shadow, we must take into account a black hole surrounded by an accretion disk. Perlick and Tsupko et al. extensively investigated the impact of plasma on a black hole shadow bib:tsupko_plasma ; bib:tsupko_deflection ; bib:tsupko_lensing ; bib:tsupko_kerr . For further insights, readers may refer to a comprehensive review on black hole shadows by Tsupko bib:tsupko_review , which also includes relevant references. The diverse characteristics of static, stationary, and dynamical black holes have been thoroughly investigated in a series of studies bib:klaudel ; bib:hod ; Virbhadra:2024pru ; Adler:2022qtb ; Virbhadra:2022iiy ; Virbhadra:2007kw ; Virbhadra:1999nm ; Virbhadra:1998dy ; bib:khoo ; bib:decaniny ; bib:shoom ; bib:sederbaum ; bib:jokhannsen ; bib:teo ; Okyay:2021nnh ; Pantig:2022gih ; Kuang:2022xjp ; Zakharov:2023lib ; Zakharov:2021gbg ; Qiao:2022hfv ; Hod:2020pim ; Lu:2019zxb ; Paithankar:2023ofw ; Aratore:2024bro ; Tsupko:2022kwi ; bib:rezzolla ; bib:rezzolla2 ; bib:expanding ; bib:vaidya1 ; bib:vaidya2 ; bib:ver_shadow ; bib:tsupko_first ; bib:understanding ; bib:joshi_shadow ; Tsukamoto:2014tja ; Tsukamoto:2017fxq ; Gomez:2024ack ; Adair:2020vso . The black hole shadow can serve as a cosmological ruler bib:tsupko_ruller ; bib:vanozi_ruller and as a means to test various black hole models bib:vanozi .
@@ III The influence of deformation on black hole shadow
We are interested in a circular light orbit, which implies $\frac{r}{d\lambda}=0$ and $\frac{d^{2}r}{d\lambda^{2}}=0$. This leads us to the following conditions on the effective potential:
@@
    $r_{ph}^{(1)}=-\frac{[r_{ph}^{(0)}-X\left(r_{ph}^{(0)}\right)]g^{\prime}\left(r_{ph}^{(0)}\right)r_{ph}^{(0)}}{\left(r_{ph}^{(0)}\right)^{2}\left(\rho_{0}+P_{0}\right)-2},$ (34)
    where $\rho_{0}$ and $P_{0}$ are the energy density and pressure of the seed metric evaluated at $r=r_{ph}^{(0)}$, respectively. The value $\left(r_{ph}^{(0)}\right)^{2}\rho_{0}$ should be less than $1$ outside the event horizon bib:wisser_dirty . Moreover, to satisfy the dominant energy condition, $|P_{0}|\leq\rho_{0}$. Thus, one can conclude:
    $\left(r_{ph}^{(0)}\right)^{2}\left(\rho_{0}+P_{0}\right)\leq 0,$ (35)

### v3

# Analyzing the Influence of Geometrical Deformation on Photon Sphere and Shadow Radius: A New Analytical Approach - Spherically Symmetric Spacetimes

@@ Abstract
In this paper, we introduce a new approach to study the behavior of the photon sphere and shadow radius. Our method uses extended gravitational decoupling and reveals two important analytic results. First, the additional matter field alters the photon sphere radius: it increases if $g^{\prime}(r_{ph}^{(0)})>0$ and decreases if $g^{\prime}(r_{ph}^{(0)})<0$ (where $g^{\prime}$ represents the derivative of a specific metric function evaluated at the original photon sphere radius). Second, the presence of the matter field can modify the black hole shadow size. If $g\left(r_{ph}^{(0)}\right)>0$, the shadow shrinks, while it grows for $g\left(r_{ph}^{(0)}\right)<0$. These findings provide a deeper insight into how matter distribution influences the characteristics of black holes and their observable features. Through a systematic framework and various illustrative examples, our investigation not only clarifies these fundamental aspects but also significantly enhances the theoretical framework of black hole astrophysics.
@@ I Introduction
One of the most convincing pieces of experimental evidence for the existence of black holes comes from the images of a black hole’s shadow, captured by the Event Horizon Telescope Collaboration bib:eht1 ; bib:eht2 ; bib:eht3 . The pioneering work on light deflection by gravitationally intense stars is attributed to Synge bib:singe . Bardeen calculated the radius of a photon sphere for a Schwarzschild black hole, demonstrating that its radius equals $3M$ bib:bardeen1 . This photon sphere manifests as a shadow, appearing as a dark spot in the observer’s sky, with a radius of $3\sqrt{3}M$ (roughly 2.5 times larger than the radius of the event horizon, $r=2M$). He also demonstrated that in the case of a rotating object, the shape of the shadow undergoes changes. However, in reality, astrophysical black holes are not isolated objects; they exist within an expanding universe and are often surrounded by luminous accretion flows, factors that significantly influence our observations. Therefore, to accurately describe a black hole shadow, we must take into account a black hole surrounded by an accretion disk. Perlick and Tsupko et al. extensively investigated the impact of plasma on a black hole shadow bib:tsupko_plasma ; bib:tsupko_deflection ; bib:tsupko_lensing ; bib:tsupko_kerr . For further insights, readers may refer to a comprehensive review on black hole shadows by Tsupko bib:tsupko_review , which also includes relevant references. The idea of using millimeter-wavelength global interferometers to image the black hole shadow at the Galactic Center was first proposed in ZAKHAROV2005479 . This prediction turned J. Bardeen’s thought experiment into a realized test of General Relativity Zakharov:2023yjl . An analytical expression for the shadow size as a function of charge was derived in Zakharov:2005ek , with the corresponding (blue) curve presented in EventHorizonTelescope:2021dqv . Additionally, the authors used this relationship to explore the existence of a black hole in higher-dimensional space 2012NewAR..56…64Z . Recent studies on M87* and Sgr A* utilized these analytical expressions for the shadow in the Reissner-Nordstrom metric to constrain the tidal charges of these black holes Zakharov:2022gwk ; Zakharov:2021gbg . While it is known that photon spheres influence black hole shadows Zakharov:1986a , a formal demonstration is needed to confirm that, for the Schwarzschild, Kerr, and Reissner-Nordstrom metrics, photons with impact parameters smaller than the critical value (corresponding to the photon sphere) are indeed captured by the black hole Zakharov:1994 . The diverse characteristics of static, stationary, and dynamical black holes have been thoroughly investigated in a series of studies Bambhaniya:2022xbz ; bib:klaudel ; bib:hod ; Virbhadra:2024pru ; Adler:2022qtb ; Virbhadra:2022iiy ; Virbhadra:2007kw ; Virbhadra:1999nm ; Virbhadra:1998dy ; bib:khoo ; bib:decaniny ; bib:shoom ; bib:sederbaum ; bib:jokhannsen ; bib:teo ; Okyay:2021nnh ; Pantig:2022gih ; Kuang:2022xjp ; Ovgun:2024zmt ; Karshiboev:2024xxx ; Uniyal:2023ahv ; Pulice:2023dqw ; Yang:2023tip ; Cimdiker:2023zdi ; Kumaran:2023brp ; Lambiase:2023hng ; Ovgun:2023ego ; Atamurotov:2022knb ; Kumaran:2022soh ; Pantig:2022qak ; Cimdiker:2021cpz ; Ovgun:2020gjz ; Ovgun:2018tua ; Zakharov:2023lib ; Qiao:2022hfv ; Hod:2020pim ; Lu:2019zxb ; Paithankar:2023ofw ; Aratore:2024bro ; Tsupko:2022kwi ; bib:rezzolla ; bib:rezzolla2 ; bib:expanding ; bib:vaidya1 ; bib:vaidya2 ; bib:ver_shadow ; bib:tsupko_first ; bib:understanding ; bib:joshi_shadow ; Tsukamoto:2014tja ; Tsukamoto:2017fxq ; Gomez:2024ack ; Adair:2020vso ; Ghosh:2023kge ; Pedrotti:2024znu ; Afrin:2022ztr ; Chen:2022nbb . The black hole shadow can serve as a cosmological ruler bib:tsupko_ruller ; bib:vanozi_ruller and as a means to test various black hole models bib:vanozi .
@@ III The influence of deformation on black hole shadow
We are interested in a circular light orbit, which implies $\frac{dr}{d\lambda}=0$ and $\frac{d^{2}r}{d\lambda^{2}}=0$. This leads us to the following conditions on the effective potential:
@@
    $r_{ph}^{(1)}=-\frac{[r_{ph}^{(0)}-X\left(r_{ph}^{(0)}\right)]g^{\prime}\left(r_{ph}^{(0)}\right)r_{ph}^{(0)}}{\left(2r_{ph}^{(0)}\right)^{2}\left(\rho_{0}+P_{0}\right)-2},$ (34)
    where $\rho_{0}$ and $P_{0}$ are the energy density and pressure of the seed metric evaluated at $r=r_{ph}^{(0)}$, respectively. The value $\left(r_{ph}^{(0)}\right)^{2}\rho_{0}$ should be less than $1$ outside the event horizon bib:wisser_dirty . Moreover, to satisfy the dominant energy condition, $|P_{0}|\leq\rho_{0}$. Thus, one can write:
    $4X^{\prime}-2=\frac{12X^{\prime\prime}}{r_{ph}^{(0)}}-10\leq 0$ (35)